OpenStudy (anonymous):

what are the x-intercepts of y=-2(x-7)(x=2)

hartnn (hartnn):

you mean \[y=-2(x-7)(x+2)\] ?? to find x-intercept, you just plug in y=0 \[-2(x-7)(x+2)=0\] can you get 2 values of 'x' from here ?

hartnn (hartnn):

by the way, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

do i use the distributive property for the -2(x-7) and then go from there??

hartnn (hartnn):

not required, if you have b(x-a) =0, you can directly take x=a. example : 4(x+10)(x-22) will give you x+10 =0, x-22 =0 x=-10,22 so what about -2(x-7)(x+2)=0 ?

OpenStudy (anonymous):

so would x=-7,2 ??

hartnn (hartnn):

see, for x-a = 0, x=a for x+c=0, x=-b so, try again with -2(x-7)(x+2)=0

hartnn (hartnn):

*for x+c=0, x=-c

OpenStudy (anonymous):

ohh ok so x=-7,-2

hartnn (hartnn):

\[(x-7)=0 \implies x=7 \\ (x+2)=0 \implies x=-2 \\ intercepts : (7,0),(-2,0)\]

OpenStudy (anonymous):

ohh ok lol i get it i just didnt see that it was a -7 i thought it was positve:P thank u!!

hartnn (hartnn):

welcome ^_^ just to clarify, if the question has x+2 ---> -2,0 if the question has x-2 ---> 2,0

OpenStudy (anonymous):

yup! and what is the point of the -2 in the begining? does it have anything to do with what they are asking?

hartnn (hartnn):

for finding x intercept, it doesn't matter. f(x) = a(x-7)(x+2) will have same intercepts, whatever a may be.

hartnn (hartnn):

same 'x' intercepts. 'y' intercepts will be different....

OpenStudy (anonymous):

ok thank u!! could u also answer another question i have on a different algebra subject?

hartnn (hartnn):

sure, ask :)

OpenStudy (anonymous):

what is the real part of the standard form of the expression (4-i)(2+3i) ??

hartnn (hartnn):

can you foil ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

do it here. do you know what 'i' is or what i=... ?

OpenStudy (anonymous):

sadly nope

hartnn (hartnn):

i is the imaginary number , i=\[\sqrt{-1}\]. so , \[i^2=... ?\]

OpenStudy (anonymous):

im not sure..

hartnn (hartnn):

\[i=\sqrt{-1} \implies i^2=-1\] now tell me what you get after foiling ?

OpenStudy (anonymous):

in the equation?

OpenStudy (anonymous):

in the equaion it would be 8+12i-2i+2i^2

hartnn (hartnn):

i'll write down steps, see if you understand. \[(4-i)(2+3i) =8+10i-3i^2 \\ but, i^2=-1 \\ =8+10i+3=11+10i\] now, real part of a+bi =a so, real part of 11+10 i=11

hartnn (hartnn):

ask if any doubts anywhere.

OpenStudy (anonymous):

were did the 10,3 and 11 come from:P

hartnn (hartnn):

8+12i-2i+2i^2---->you didn't foil correctly 8+12i-2i-3i^2 = 8+ 10 i -3(-1) got this ?

hartnn (hartnn):

-i * +3i = -3i^2

OpenStudy (anonymous):

ohh yup i got it

hartnn (hartnn):

now any more doubts on how real part = 11 ?

OpenStudy (anonymous):

nope i got it!

hartnn (hartnn):

good to hear , glad to help :)

OpenStudy (anonymous):

thank u! i really like this website! im glad i found it!!